OpenStudy (anonymous):

Assume you have 6 dimes and 5 quarters (all distinct), and you select 5 coins. Total ways selection can be made= 462 How many ways can the selection be made so that at least 4 coins are dimes?______________

OpenStudy (anonymous):

(6C4)+(7C1)

OpenStudy (anonymous):

6x5x4x3 -------- = 15 +7= 22 4x3x2x1 Is that right?

OpenStudy (anonymous):

that didnt work

OpenStudy (anonymous):

6*5*4*3*2*1 7*6*5*4*3*2*1 ------------ + -------------- 4*3*2*1*2*1 1*6*5*4*3*2*1

Directrix (directrix):

Two separate cases: That would be the sum of the numbers of ways to choose exactly 4 dimes and then exactly 5 dimes.. C(6,4) times C(5,1) + C(6,5) times C(5, 0) = 81.

Directrix (directrix):

I included the case of 6 dimes and messed up the first time.

OpenStudy (anonymous):

That did it direct.... awesome man

OpenStudy (anonymous):

it is hard for me to id the cases in a lot of these

Directrix (directrix):

It is tricky.

OpenStudy (anonymous):

1 more i couldn't do from that section i am going to post.... thank you so much

Directrix (directrix):

Ok, I'll look for it.